A line with a slope of 1/2 is drawn thru point P at (-2,,2). Moving from left to right, what is the next ordered pair of whole numbers on the line?
.............................................

How does one figure out this answer?

Answer by abhinav75(8)

The slope of the line is 1/2.
now, let us take the equation of the line as:
y = mx + c
now, m = 1/2
so, y = 1/2x + c
=> 2y = x + 2c
now, it passes through (-2,2)
2(2) = -2 + 2c
=> 4 = -2 + 2c
=> 2c = 6
=> c= 3
so, the equation is : y = 2x + 3
to find next ordered pair..
now, see that whenever x is a whole number, y wud be a whole number as well.
so, if we take x = 0, y =3.
so, the next pair wud be (0,3)
Please note that whole numbers start from 0.